OpenStudy (anonymous):

i did this last year, but can someone explain to me how to add and subtract radicals? example below.

OpenStudy (anonymous):

\[\sqrt{2}-\sqrt{18}+\sqrt{6}\]

OpenStudy (precal):

like terms

OpenStudy (precal):

make them like terms if possible

OpenStudy (anonymous):

my mathbook say the answer is \[-2\sqrt{2}+2\sqrt{6}\], but im not sure how to get there. yeah, like terms but i still don't understand.

OpenStudy (anonymous):

also algebra was last year so i feel silly :/

OpenStudy (precal):